OpenStudy (anonymous):

in need of help with some equations 64=74-x

OpenStudy (anonymous):

x=74-64=10

OpenStudy (anonymous):

so x=10 ?

OpenStudy (anonymous):

@sambhav__jain

OpenStudy (anonymous):

Write an equation to represent this verbal statement: three times the sum of n and 7 is 15 can anyone help ?

OpenStudy (anonymous):

3(7+n)=15

OpenStudy (anonymous):

A camp director has 255 campers who need to divided into 15 equal groups, each containing an unknown number of campers How can i write this as an equation ?

OpenStudy (anonymous):

x=255/15=17

OpenStudy (anonymous):

medal pls

OpenStudy (anonymous):

Sherice has 16 bags, each containing the same number of chewy candies. She also has eight jars, and each jar has the same number of chewy candies. A jar of candies has twice as many candies as a bag of candies. Sherice gives away six bags and then hands her friend Klara 10 candies. In the end, Sherice is left with 302 chewy candies. Each bag has how many chewies in it ?